# GQE: Generalized Query Expansion for Enhanced Text-Video Retrieval

### v1

@@ 4.2.1 Query Expansion of Training Set
4.2.2 Query Expansion of Testing
@@
We use 1 original query and 10 expanded queries on the MSR-VTT dataset to evaluate and compare different types of query generation and result aggregation methods.
@@
Query Generation. We first compare the two types of query generation methods. For the NLTK-based method, we utilize the NLTK [6] tool to identify representative terms (nouns and verbs) in the query and generate new queries by replacing these terms with their synonyms obtained from the NLTK tool. For the LLM-based method, we instruct the LLM (GPT-4) to generate queries through specific prompts. As shown in Tab. 6, under all the result aggregation methods, LLM always outperforms the NLTK-based method. Notably, the upper bound results, which remove the influence of aggregation methods, provide a clearer distinction between the two query generation methods. This performance gap demonstrates the effectiveness of LLM as a tool for query expansion in text-video retrieval. Despite being slightly less effective than LLM, it’s worth noting that NLTK still achieves significant improvement compared with baseline. It should be acknowledged that NLTK-based methods can also be a cheaper option with some compromise on the performance.
@@
Result Aggregation. Next, we investigate different methods for aggregating the retrieval results from multiple queries. The results in Tab. 6 show that certain methods, such as concatenating queries and mean-pooling of text embeddings, can lead to even worse performance than the baseline. Among these methods, majority voting is a simple yet effective one, outperforming other methods. The most interesting part is the Oracle result. Firstly, it shows the huge potential of testing query expansion with the significant performance improvement. Secondly, the gap between the Oracle result and other methods reveals that there are still a large room for improvement. Finally, the Oracle result is achieved without compromising on computational cost while majority voting with all the 11 queries appears to be inefficient. These findings inspire us to develop a dedicated query expansion module to further improve retrieval performance as well as reduce computational cost.
Query Selection aims to enhance retrieval performance to approach the Oracle and improve inference efficiency. In Tab. 7, we report the R@1 retrieval performance under various query selection settings. Firstly, we find that randomly selecting a specific number of queries can result in promising performance. The primary drawback of random query selection is its lack of stability, a characteristic that should be avoided in retrieval systems. Next, we adapt the determinantal point processes (DPPs) [29] into the query selection task. In particular, we utilize MAP inference of DPPs introduced in [10] for fast and deterministic inference. It returns a subset of points of a pre-specified size (i.e. $k$) that are maximally diverse. The results show that $k$-DPP can be a viable implementation of query selection. Finally, we observe that the proposed Farthest Query Sampling algorithm achieves competitive or better performance in a deterministic way.
Compared to a single query or random query sampling, FQS is able to cover a larger area in the embedding space, naturally increasing the diversity. When comparing $k$-DPP and FQS, the main difference is that $k$-DDP selects queries from the list of candidates comprised of GT query and expanded queries. In other words, $k$-DDP treat GT query and expanded queries equally. It is possible that the GT query will not be selected. In contrast, our FQS always select the GT query as the anchor then select other diverse queries from the expanded queries. The involvement of GT query essentially ensures the relevance of text query and establishes a solid retrieval lower bound, allowing FQS to achieve better retrieval performance. We use the text encoder in our retrieval model for embedding extraction in the experiments unless specified.
@@
4.3 Discussion
We provide more results in the appendix, including implementation details, experiments on model architectures rather than X-Pool [23] showing that our data-centric approach is compatible with model-centric approaches; discussion about the interesting oracle query selection; video-to-text retrieval result; and qualitative examples that demonstrate how the expanded queries aid in text-video retrieval. We also thoroughly discuss the limitation and future work.
@@ 5 Conclusion
In this work, we address the one-to-many problem in text-video retrieval task resulting from information imbalance between video and text queries. We introduce GQE, a data-centric approach to address this issue in both training and testing phases. For training dataset, we expand text queries through comprehensive video captioning, ensuring full video scene coverage. During testing, we use large language models to generate multiple relevant yet diverse text queries to enhance retrieval performance. Moreover, we apply farthest query sampling for effective query selection, improving retrieval performance and reducing computational costs. Extensive experiments prove GQE has consistent improvements over four benchmarks, surpassing current state-of-the-art methods. Our work serves as an inspiration for future research efforts, encompassing both model-centric methods and data-centric enhancements.
@@ A.1 Experiment Setups and Implementation Details
Datasets. MSR-VTT [66] contains a total of 10K video clips, each having 20 captions. We utilize the Training-9k subset for training and the test-1K-A subset for evaluation. MSVD [64] contains 1,970 videos with 80K captions, with 40 captions on average per video. There are 1200, 100, and 670 videos in the train, validation, and test sets, respectively. LSMDC [49] consists of 118,081 video clips sourced from 202 movies with one caption corresponding to each clip. Evaluation is conducted on a test set of 1,000 videos from movies disjoint from the train and validation sets. VATEX [61] collects around 35K videos with multiple text annotations in both English and Chinese for each video. There are around 26K videos for training, 1,500 for validation, and 1,500 for testing.
@@
Implementation Details. We initialize our backbone from the pre-trained CLIP ViT-B/32. Following previous work [23], the text-conditioned pooling module is randomly initialized. Our models undergo end-to-end training on each dataset. For the training hyper-parameters, we also follow X-Pool [23]. The model is trained for 3 epochs on the expanded training set and 5 epochs on the standard training set. A cosine scheduler [37] is employed to decay the learning rate. For all experiments, we follow previous works by uniformly sampling 12 frames from each video and resizing them to 224x224. For query expansion of training set, we use blip2-opt-2.7b-coco as the captioner. For query expansion during testing, we utilize GPT-4 model through the API  [^2] https://openai.com/blog/openai-api to generate queries. We train our model using NVIDIA A10 24 GPU. The training takes around 12 hours.
A.2 Experiment on Vanilla Model
@@
In our main paper, we utilize the pre-trained CLIP and the text-conditioned pooling module as the retrieval model. The text-conditioned pooling is designed to capture video frames relevant to the text query, effectively alleviating the information imbalance. In this section, we demonstrate that our data-centric approach is not coupled with a specific architecture design, such as the text-conditioned pooling. We conduct experiments on the vanilla CLIP model with a mean-pooling strategy to obtain the video embedding. The results are reported in Tab. 8.
@@
Table 8: Text-to-video retrieval results with vanilla CLIP mean-pooling model on MSRVTT [65].
# GQE Training GQE Testing R@1 R@5 R@10 MdR MnR
@@
The comparison between Row 1 and Row 2, as well as between Row 3 and Row 4, illustrate the effectiveness of GQE Training. Without GQE Testing (Row 1 vs. Row 2), GQE Training achieves an improvement of 0.7% on R@1. The performance gain becomes more obvious when combined with GQE Testing (+2.3% on R@1, Row 3 vs. Row 4). Finally, the 6.6% performance gap between the results in Row 1 and Row 4 further validates the benefits of our full approach.
A.3 GQE Testing as a Training-free Tool
In this section, we demonstrate that the proposed query expansion of testing can be applied to a wide range of text-video retrieval models as a training-free tool for performance improvement. We conduct evaluations on the MSR-VTT dataset [65] using four widely recognized methods. 1) For CLIP [47] zero-shot, we utilize the pre-trained CLIP model and employ mean-pooling to obtain the video embedding in a zero-shot manner; 2) CLIP4Clip [38], a classical method in the text-video retrieval domain; 3) X-Pool [23], adopted as the base model in our work; 4) CLIP-VIP [68], recognized as one of the current state-of-the-art methods. To maintain consistency, all experiments utilize the ViT-B/32 as the retrieval backbone and CLIP text encoder for query selection. The baseline results are reproduced from their released codebases.
Table 9: Text-to-video retrieval results with various methods on the MSR-VTT [65]. The result are reproduced based the released codebase of each method respectively.

### v2

@@ 4.2.1 Text Enrichment of Training Phase
4.2.2 Text Enrichment of Retrieval Phase
@@
We use 1 original query and 10 enriched queries on the MSR-VTT test set to evaluate and compare different types of query generation and result aggregation methods.
@@
Query Generation. We first compare the two types of query generation methods. For the NLTK-based method, we utilize the NLTK Bird et al. (2009) tool to identify representative terms (nouns and verbs) in the query and generate new queries by replacing these terms with their synonyms obtained from the NLTK tool. This strategy can be regarded as traditional query expansion. For the LLM-based method, we instruct the LLM (GPT-4) to generate text descriptions by reasoning the visual scene. As shown in Tab. 6, under all the result aggregation methods, LLM always outperforms the NLTK-based method. Notably, the oracle result, which is achieve by using one single text query and removes the influence of aggregation methods, provide a clearer distinction between the two query generation methods. We will elaborate the detail of how we achieve the oracle performance in Appendix. This performance gap demonstrates the effectiveness of using LLM as a tool for text query enrichment, thanks to its exceptional capabilities on high-level language understanding, reasoning, and generation.
@@
Result Aggregation. Next, we investigate different methods for aggregating the retrieval results from multiple queries. The results in Tab. 6 show that certain methods, such as concatenating queries and mean-pooling of text embeddings, can lead to even worse performance than the baseline. Among these methods, majority voting is a simple yet effective one, outperforming other methods. The most interesting part is the Oracle result. Firstly, it shows the huge potential of text enrichment in retrieval phase with the significant performance improvement. Secondly, the gap between the Oracle result and other methods reveals that there are still a large room for improvement. Finally, the Oracle result is achieved without compromising on computational cost, while majority voting with all the 11 queries appears to be inefficient. These findings inspire us to develop a dedicated query selection module to further improve retrieval performance as well as reduce computational cost.
4.2.3 Impact of Query Selection Mechanism
Query selection aims to enhance retrieval performance to approach the Oracle performance and improve inference efficiency. In Tab. 7, we report the R@1 retrieval performance under various query selection strategies and selection number $k$. Firstly, we employ a random selection strategy on the baseline model and run the retrieval with 4 different random seeds, resulting the first 4 rows of the table. It can be observed that randomly selecting a specific number of queries can lead to promising performance. For example, when $k=2$, the R@1 accuracy can achieve $49.6$ or even $50.2$. The performance variation among different $k$ indicates that not all the enriched queries contribute equally.
The primary drawback of random query selection is its lack of stability, making the retrieval result of each round different. This characteristic should be avoided in retrieval systems. We demonstrate that the our proposed Farthest Query Sampling achieves competitive or better performance under the same setting of $k$. The performance improvement can be attributed to that our query selection method considers both relevance and diversity of text queries, being able to cover a larger area in the embedding space and naturally increasing successful retrieval rate.
We also compare our FQS with another deterministic method, named determinantal point processes (DPPs) Kulesza et al. (2012). In particular, we utilize MAP inference of DPPs introduced in Chen et al. (2018) for fast and deterministic inference. DPP is similar to our FQS in implementation. It returns a subset of points of a pre-specified size (i.e. $k$) that are maximally diverse. The results show that $k$-DPP can be a viable implementation of query selection. However, it exhibits a slightly worse performance than FQS. We analyze that the main difference lies in the initial text query. In FQS, we always take the user query as the starting point and probe enriched queries to maximize the diversity. The involvement of ground-truth user query essentially ensures the relevance of text query and establishes a solid retrieval lower bound, allowing FQS to achieve better retrieval performance. In contrast, $k$-DPP treat the ground-truth user query and the enriched queries equally. It maximizes the diversity aspect from a global view, while neglecting the relevance aspect. Thus, it is possible the user query would not be selected. making the retrieval be sensitive to some noisy enriched queries.
@@ 4.2.3 Impact of Query Selection Mechanism
4.2.4 Impact of LLM Choice in Retrieval Phase
Table 8: Experiments conducted on MSR-VTT Xu et al. (2016b) dataset, assessing the impact of LLM employed in retrieval text query enrichment.
LLM # Param. Text Query R@1 R@5 R@10
- - GT User Query 47.7 71.9 82.1
GPT-4 Unknown Farthest Enriched Query 41.6 66.2 76.3
Phi-3.5 3.8B Farthest Enriched Query 36.7 62.1 73.7
GPT-4 Unknown FQS ($k$=2) 52.1 76.8 86.3
Phi-3.5 3.8B FQS ($k$=2) 51.5 77.7 86.5
In the retrieval phase, the text queries are enriched by an LLM, which plays a crucial role. Throughout the paper, we use GPT-4 API to achieve the text enrichment. In this section, we use an open-source, light-weight LLM, Phi-3.5 Abdin et al. (2024) to do the text enrichment task in retrieval phase. The experiment results are shown in Table 8.
Firstly, when using the farthest one of the enriched queries, we observe that both GPT-4 and Phi-3.5 exhibit inferior retrieval performance compared to using GT user query. The performance gap indicates that naively using LLM to enrich the text query would introduce noise. In the two LLMs, Phi-3.5 shows much worse performance than GPT-4. This aligns well with our expectation, as the Phi-3.5 model is known to be weaker than GPT-4, introducing more noise in the enriched text queries.
Next, we apply our proposed Farthest Query Sample algorithm to the enriched text queries and conduct retrieval. Surprisingly, we observe only a small performance gap between the two LLMs, 0.5% in Rank-1. This experiment proves that our method is robust to the LLM. The FQS algorithm and the majority voting aggregation can mitigate large noise. Moreover, the exceptional performance of Phi-3.5 unveil of opportunity of using small language model, further improving inference efficiency.
4.3 More Results
We provide more results in the appendix, including A.1: Experiment setups and implementation details; A.3: Details and more interesting findings of the oracle query selection; A.4: Analysis of computational efficiency; A.5 and A.6: experiments on model architectures beyond X-Pool showing that our data-centric approach is compatible with a wide range of model-centric approaches; A.7: prompt in LLM;  A.9: qualitative examples that demonstrate how the enriched text queries in retrieval aid in text-video retrieval. We also thoroughly discuss the limitation and future work.
@@ 5 Conclusion
In this work, we address the information asymmetry problem between video and text in the text-video retrieval task. We take a novel data-centric perspective to address this issue at multiple stages. During training, we enrich text descriptions through an event-aware, comprehensive video captioning approach, ensuring full video scene coverage. In the retrieval process, we use large language models to enrich the text queries by reasoning the semantic concepts. We introduce the concept of Oracle Query to reveal the significant potential of data-centric investigation in retrieval. Moreover, we design a farthest query sampling algorithm for effective query selection, improving retrieval performance and reducing computational costs. Extensive experiments prove that our method has consistent improvements over four benchmarks, surpassing current state-of-the-art methods. Our work serves as an inspiration for future research efforts, encompassing both model-centric methods and data-centric enhancements.
Acknowledgement
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-030).
Ethics Statement
This research does not involve human subjects or the collection of personally identifiable information. The datasets used in this work are publicly available and have been previously vetted for ethical use in academic research. We ensured compliance with relevant data licensing agreements and conducted all experiments in accordance with the terms of use of the datasets. The proposed framework that involves the usage of a large language model (LLM), was designed to improve the fairness and accuracy of cross-modal retrieval systems. However, we acknowledge that LLMs may inadvertently encode and propagate biases present in the training data. To mitigate this, we adopted a data-centric approach aimed at generating diverse and representative queries. We encourage future work to continue monitoring and addressing bias in these models to prevent unintended societal harm. Additionally, our method was designed with computational efficiency in mind to reduce resource consumption and ensure a more environmentally sustainable research practice. No conflicts of interest or external sponsorship influenced the outcomes of this research.
Reproducibility Statement
We have made every effort to ensure that our results are fully reproducible. Detailed descriptions of the proposed framework, including the data processing pipeline, model architecture, and training procedures, are provided in the main paper as well as the comprehensive appendix. To facilitate reproducibility, all datasets used in our experiments are publicly datasets, and the specific processing steps are sufficiently described. We will make the code, data, and pre-trained model public.
@@ A.1 Experiment Setups and Implementation Details
Datasets. MSR-VTT Xu et al. (2016a) contains a total of 10K video clips, each having 20 captions. We utilize the Training-9k subset for training and the test-1K-A subset for evaluation. MSVD Wu et al. (2017a) contains 1,970 videos with 80K captions, with 40 captions on average per video. There are 1200, 100, and 670 videos in the train, validation, and test sets, respectively. LSMDC Rohrbach et al. (2015a) consists of 118,081 video clips sourced from 202 movies with one caption corresponding to each clip. Evaluation is conducted on a test set of 1,000 videos from movies disjoint from the train and validation sets. VATEX Wang et al. (2019) collects around 35K videos with multiple text annotations in both English and Chinese for each video. There are around 26K videos for training, 1,500 for validation, and 1,500 for testing.
@@
Implementation Details. We initialize our backbone from the pre-trained CLIP ViT-B/32. Following previous work Gorti et al. (2022), the text-conditioned pooling module is randomly initialized. Our models undergo end-to-end training on each dataset. For the training hyper-parameters, we also follow X-Pool Gorti et al. (2022). The model is trained for 3 epochs on the enriched training set and 5 epochs on the standard training set. A cosine scheduler Loshchilov & Hutter (2016) is employed to decay the learning rate. For all experiments, we follow previous works by uniformly sampling 12 frames from each video and resizing them to 224x224. For text enrichment in training, we use blip2-opt-2.7b-coco as the captioner. For text enrichment in retrieval, we utilize GPT-4 model through the API  [^1] https://openai.com/blog/openai-api to generate queries. We train our model using NVIDIA A10 24G GPU. The training takes around 12 hours.
A.2 Additional Ablation of FQS
Table 9: Comparing FQS with other variants.
Query Selection (k=2) Rank-1 Rank-5 Rank-10
Random 49.6 76.1 85.9
Nearest Neighbor 50.4 76.0 85.7
Farthest Neighbor 51.4 76.5 85.8
k-DPP 50.9 76.7 85.8
FQS (s=0.5) 52.1 76.8 86.3
FQS (s=0.75) 52.1 76.8 86.3
FQS (s=0.85) 51.6 76.7 86.2
FQS (s=0.95) 49.8 76.1 86.3
NQS 49.9 75.8 86.1
FQS 52.1 76.8 86.3
The Farthest Query Sampling (FQS) algorithm proposed in the main paper plays an important role on selecting relevant yet diverse queries. In this section, we compare this algorithm with a wide range of settings to further demonstrate its effectiveness. The results are shown in Tab. 9.
  • Random serves as a baseline, which randomly select $k$ queries from the enriched queries.
  • Nearest Neighbor selects $k$ queries that are closest to the user query from the enriched queries.
  • Farthest Neighbor selects $k$ queries that are farthest to the user query from the enriched queries.
  • $k$-DPP is classical algorithm selecting a subset with maximal diversity from a set of data points.
  • FQS(s=) is an variant of FQS suggested by one reviewer during the review process of the paper. The value of “s" is a minimum similarity threshold. In implementation, when iteratively selecting queries in FQS, we add an additional condition that requires the selected queries to meet the minimum similarity. If no query meets this condition, we will use the original user query instead.
  • NQS is another variant of FQS suggested by the reviewer, denoting Nearest Query Sampling, which minimizes the minimum distance between selected queries.
The comparison in Tab. 9 can further verify the effectiveness of the proposed FQS algorithm.
A.3 Understanding Oracle Query Selection
The oracle query selection is implemented as follows: for one specific instance, among 1 GT query and $k$ enriched queries, we always select the query that yields the best ranking, where the best ranking refers to that the target video is ranked highest. If there are multiple such queries, we prioritize the GT query. We call it ‘Oracle’ because we need to access the ground-truth matching during the query selection, which is impractical in real application. Following the rule above, we stats the percentage of GT query vs. enriched query in the selected oracle query. The statistic is conducted on $1,000$ test samples of MSR-VTT dataset, $k=10$. The result shows that $59.5\%$ of selected queries come from the GT query set, while $40.5\%$ of selected queries are from the enriched queries. This actually provides an explanation of the huge performance gap between baseline and oracle, suggesting that the text queries can be specially investigated to enhance the retrieval.
A.4 Analysis of computational cost
Table 10: Analysis of computational cost.
Model Term $N_{t}$ $N_{e}$ $N_{v}$ Memory Infer. time
Txt Enc. $C_{txt\_enc}$ 1 0 - 602 MB 6.30 ms
Txt Enc. (Text Enrich.) $(1+N_{e})\otimes C_{txt\_enc}$ 1 10 - 618 MB 8.65 ms
Similarity $C_{sim}$ 1 0 100,000 10096 MB 96.71 ms
Similarity (Text Enrich.) $(1+N_{e})\otimes N_{v}\otimes C_{sim}$ 1 10 100,000 16258 MB 233.80 ms
Similarity (Query Select.) $(1+N_{e})\otimes N_{v}\otimes C_{sim}$ 1 2 100,000 11490 MB 139.27 ms
Full Infer. (Baseline) $C_{total}-C_{LLM}$ 1 0 100,000 - 103.01 ms
Full Infer. (Text Enrich.) $C_{total}-C_{LLM}$ 1 10 100,000 - 242.45 ms
Full Infe. (Query Select.) $C_{total}-C_{LLM}$ 1 2 100,000 - 147.92 ms
We decompose the computational cost of a CLIP-based retrieval model into several parts. Formally, given $N_{t}$ text queries and $N_{v}$ candidate videos, supposing each text query is enriched with $N_{e}$ additional text queries, the computational cost can be expressed as follows:
$\begin{split}C_{total}=(N_{t}+N_{t}\times N_{e})\otimes C_{txt\_enc}+N_{v}\otimes C_{vid\_enc}+\\ (N_{t}+N_{t}\times N_{e})\otimes N_{v}\otimes C_{sim}+N_{t}\otimes C_{LLM},\end{split}$ (10)
where $C_{txt\_enc}$ and $C_{vid\_enc}$ are the inference cost of the text encoder and video encoder, respectively. $C_{sim}$ is the cost of computing the cross-modal similarity. $C_{LLM}$ is the inference cost of LLM. The operation of $\otimes$ indicates that the the cost $C$ is influenced by the factors $N$. Note that $\otimes$ is different from traditional multiplication operation $\times$, as it can be parallelized and accelerated by GPU.
In practical applications, we usually use one query to search among a large volume of candidate videos. So, we first set $N_{t}=1$ for simplicity. Besides, the video embeddings are usually pre-extracted and stored in a database. Thus, the $C_{vid\_enc}$ can be omitted. For each query, the total cost can be simplified as:
$C_{total}=(1+N_{e})\otimes C_{txt\_enc}+(1+N_{e})\otimes N_{v}\otimes C_{sim}+C_{LLM}.$ (11)
Based on this, we suggest two ways to reduce the computational cost. First, our query selection mechanism can effectively reduce $N_{e}$ without sacrificing performance. In fact, it even further boost the performance by a large margin. It is worth mentioning that in real world applications, $N_{v}$ can be super large. Thus, reducing $N_{e}$ would be effective on improving the overall efficiency. Second, our approach exhibits strong robustness to the selection of LLM, as discussed in Sec 4.2.4 of the paper. It implies that we are allowed to employ lightweight LLM to reduce $C_{LLM}$ and improve efficiency.
To help understand the inference cost more intuitively, we provide an example in Table 10. In a practical retrieval scenario, we may have one user query and a large volume of candidate videos, say $N_{t}=1$ and $N_{v}=100,000$. The inference time are computed by averaging 1,000 randomly generated examples. We observe that, with the help of GPU parallelization, increasing $N_{e}$ does not lead to significant inference burden in text encoder. However, the similarity computation cost a long inference time. This is because X-Pool requires a text-conditioned pooling, as introduced in the main paper. Finally, we show that vanilla text enrichment strategy increase the per query inference time from $103.01ms$ into $242.45ms$. With the help of query selection, the inference time decrease to $147.92ms$, validating the effectiveness of query selection. Not that this inference time does not take $C_{LLM}$ into account, as it varies depends on different LLMs. We acknowledgement that the inference cost of LLM could be large. However, it can be reduced by selecting lightweight model or more advanced techniques, e.g., KV Cache Pope et al. (2023). Our main goal is to unveil the power of data in text-video retrieval. We leave this optimization as a future work.
A.5 Experiment on Vanilla Model
@@
In our main paper, we utilize the pre-trained CLIP and the text-conditioned pooling module as the retrieval model. The text-conditioned pooling is designed to capture video frames relevant to the text query, effectively alleviating the information asymmetry. In this section, we demonstrate that our data-centric approach is not coupled with a specific architecture design, such as the text-conditioned pooling. We conduct experiments on the vanilla CLIP model with a mean-pooling strategy to obtain the video embedding. The results are reported in Tab. 11.
@@
Table 11: Text-to-video retrieval results with vanilla CLIP mean-pooling model on MSRVTT Xu et al. (2016b).
# Text Enrich. Training Text Enrich. Retrieval R@1 R@5 R@10 MdR MnR
@@
The comparison between Row 1 and Row 2, as well as between Row 3 and Row 4, illustrate the effectiveness of Text Enrichment in Training. In the baseline model (Row 1 vs. Row 2), Text Enrichment in Training achieves an improvement of 0.7% on R@1. The performance gain becomes more obvious when combined with Text Enrichment in Retrieval (+2.3% on R@1, Row 3 vs. Row 4). Finally, the 6.6% performance gap between the results in Row 1 and Row 4 further validates the benefits of our unified framework.
A.6 Text Enrichment in Retrieval as a Training-free Tool
In this section, we demonstrate that the proposed Text Enrichment in Retrieval can be applied to a wide range of text-video retrieval models as a training-free tool for performance improvement. We conduct evaluations on the MSR-VTT dataset Xu et al. (2016b) using four widely recognized methods. 1) For CLIP Radford et al. (2021) zero-shot, we utilize the pre-trained CLIP model and employ mean-pooling to obtain the video embedding in a zero-shot manner; 2) CLIP4Clip Luo et al. (2022), a classical method in the text-video retrieval domain; 3) X-Pool Gorti et al. (2022), adopted as the base model in our work; 4) CLIP-VIP Xue et al. (2022), recognized as one of the current state-of-the-art methods. To maintain consistency, all experiments utilize the ViT-B/32 as the retrieval backbone and CLIP text encoder for query selection. The baseline results are reproduced from their released codebases.
Table 12: Text-to-video retrieval results with various methods on the MSR-VTT Xu et al. (2016b). The result are reproduced based the released codebase of each method respectively.